# The reflection complexity of sequences over finite alphabets

### v1

@@ 2 General results
Let $\mathbf{x}$ be a primitive morphic sequence. We know that $\mathbf{x}$ is uniformly recurrent. Thus, from Theorem 11, $\mathbf{x}$ is either reversal-closed, or else it has no long palindromes. If $\mathbf{x}$ is reversal-closed, then, by Corollary 21, we have $r_{\mathbf{x}}(n)\sim\frac{1}{2}\rho_{\mathbf{x}}(n)$. Otherwise, $\mathbf{x}$ has no long palindromes, then, still from Theorem 11, we have that $r_{\mathbf{x}}(n)=\rho_{\mathbf{x}}(n)$ for $n$ large enough. ∎
@@ 3 Eventually periodic sequences
From Theorem 22 both sequences $(r_{\mathbf{x}}(2n))_{n\geq 0}$ and $(r_{\mathbf{x}}(2n+1))_{n\geq 0}$ are non-decreasing. Also, from the inequality $\frac{1}{2}\rho_{\mathbf{x}}(n)\leq r_{\mathbf{x}}(n)\leq\rho_{\mathbf{x}}(n)$ in Theorem 9, and the fact that the sequence $(\rho_{\mathbf{x}}(n))_{n\geq 0}$ is non-decreasing, we have that either the three integer sequences $(r_{\mathbf{x}}(2n))_{n\geq 0}$, $(r_{\mathbf{x}}(2n+1))_{n\geq 0}$, and $(\rho_{\mathbf{x}}(n))_{n\geq 0}$ are all bounded, or else none of them is. Furthermore, we know that $(\rho_{\mathbf{x}}(n))_{n\geq 0}$ is bounded if and only if the sequence $\mathbf{x}$ is eventually periodic (Theorem 5(d) above). Hence, we have two cases depending on the periodicity of $\mathbf{x}$.
@@
    If $\mathbf{x}$ is not eventually periodic, its factor complexity is not bounded: thus both sequences $(r_{\mathbf{x}}(2n))_{n\geq 0}$ and $(r_{\mathbf{x}}(2n+1))_{n\geq 0}$ tend to infinity.
@@
Remark 34.
Theorem 35, Corollary 36 and Remark 38 below give more precise results for the growth of the reflection complexity of eventually periodic and non-eventually periodic sequences.
@@ 4 Sturmian sequences and generalizations
In this section, we study Sturmian sequences as well as some generalizations.
4.1 Sturmian sequences
First we state the following result, which notably characterizes Sturmian sequences in terms of their reflection complexity.
@@ 4.1 Sturmian sequences
Let $\mathbf{x}$ be a non-eventually periodic sequence on a finite alphabet.
@@
and $G_{n}$ be the subgroup of ${\mathcal{S}}_{n}$ generated by $\sigma_{n}$, i.e., the group $\{\sigma_{n},{\rm id}_{n}\}$. The number of distinct orbits of $\{1,2,\ldots,n\}$ under $G_{n}$ is equal to $n/2$ if $n$ is even, and to $(n+1)/2$ if $n$ is odd, which can be written $\lfloor(n+1)/2\rfloor$ in both cases. Thus, applying [33, Theorem 1] proves the first item of the theorem and the implication $\Longrightarrow$ of the second item.
To prove the last assertion, suppose that $\mathbf{x}$ is a Sturmian sequence. We know that any Sturmian sequence is reversal-closed (see [43, Theorem 4, p. 77], where reversals are called mirror images). Furthermore, it is proved in [43, Theorem 5, p. 77] that a sequence is Sturmian if and only if it has one palindrome of all even lengths and two palindromes of all odd lengths. Now, from Theorem 9(b) we have that
@@
This ends the proof. ∎
With regard to the above referenced work of Charlier et al. [33]; also see the related and recent work by Luchinin and Puzynina [61].
@@
Remark 37.
Actually, it is possible to prove the first part of the proof of Corollary 36 without using Theorem 35 in the case where the integer $n$ is even. Namely, Theorem 9 implies that $\rho_{\mathbf{x}}(n)/2\leq r_{\mathbf{x}}(n)$ for all $n$. So that, if $r_{\mathbf{x}}(n_{0})\leq\lfloor\frac{n_{0}+1}{2}\rfloor$ for some $n_{0}$, then $\rho_{\mathbf{x}}(n_{0})\leq 2\lfloor\frac{n_{0}+1}{2}\rfloor=n_{0}$ if $n_{0}$ is even. Then we apply the Morse–Hedlund theorem (Theorem 5).
Remark 38.
As in Remark 6, there is another “growth gap”. Namely, an easy consequence of Theorem 9 above is that, for any sequence $\mathbf{x}$ that is neither eventually periodic, nor quasi-Sturmian, one has $r_{\mathbf{x}}(n)-\frac{n}{2}\to+\infty$ when $n\to+\infty$ (recall the definition of quasi-Sturmian in Remark 6; and use the fact that, for a sequence $\mathbf{x}$ that is neither eventually periodic nor quasi-Sturmian, one has that $\rho_{\mathbf{x}}(n)-n\to+\infty$). So that we can have the following possibilities for a sequence $\mathbf{x}$:
@@
    $r_{\mathbf{x}}(n)$ is bounded, which happens if and only if $\mathbf{x}$ is eventually periodic. Then both sequences $(r_{\mathbf{x}}(2n))_{n}$ and $(r_{\mathbf{x}}(2n+1))_{n}$ are eventually constant;
@@
    $r_{\mathbf{x}}(n)=1+\lfloor\frac{n+1}{2}\rfloor$ for all $n$, which happens if and only if $\mathbf{x}$ is Sturmian. Note that $r_{\mathbf{x}}(n)-\frac{n}{2}>\frac{1}{2}$ for any non-eventually periodic sequence, and that $r_{\mathbf{x}}(n)-\frac{n}{2}\in\{1,\frac{3}{2}\}$ for any Sturmian sequence;
  • (c)
    $r_{\mathbf{x}}(n)-\frac{n}{2}$ is bounded, and $\mathbf{x}$ is not Sturmian. This implies that $\rho_{\mathbf{x}}(n)-n$ is bounded (use Theorem 9), hence that $\mathbf{x}$ is quasi-Sturmian;
  • (d)
    $r_{\mathbf{x}}(n)-\frac{n}{2}$ is not bounded and $\mathbf{x}$ is quasi-Sturmian.
  • (e)
    $r_{\mathbf{x}}(n)-\frac{n}{2}$ tends to infinity. This is the case where $\mathbf{x}$ is neither eventually periodic nor quasi-Sturmian.
Note that both behaviors $r_{\mathbf{x}}(n)-\frac{n}{2}$ bounded (Item (c)) or $r_{\mathbf{x}}(n)-\frac{n}{2}$ not bounded (Item (d)) are possible for quasi-Sturmian sequences. Namely, if we start from the binary Fibonacci sequence $\mathbf{f}$ (fixed point of the morphism $0\to 01$, $1\to 0$), and apply two particular morphisms, then (using Walnut), we have for the corresponding quasi-Sturmian sequences:
  • (a)
    the image of $\mathbf{f}$ under the morphism $0\to 0101$, $111\to 1$ $1$ is reversal closed and has arbitrarily long palindromes; its reflection complexity is given by $r(n)-\frac{n}{2}=\frac{9}{2}$ for $n\geq 7$ odd and $r(n)-\frac{n}{2}=3$ for $n\geq 8$ even;
  • (b)
    the image of $\mathbf{f}$ under the morphism $0\to 01101$, $1\to 10100$ is not reversal-closed, and has no large palindrome; its complexity has the property that $r(n)=n+9$ for $n\geq 11$.
It is also worth noting that in [38, p. 133], it is indicated that, for any Sturmian sequence $\mathbf{x}$ with values in $\{0,1\}$, the image of $\mathbf{x}$ by the morphism $f$ defined by $f(0)=011001$, $f(1)=001011$ is a quasi-Sturmian sequence without long palindromes (the authors call such a sequence a non-palindromic sequence, using the terminology of [53]). More precise results on the reflection complexity of quasi-Sturmian sequences are given in the next subsection.
4.2 Quasi-Sturmian sequences
Recall that any quasi-Sturmian sequence $\mathbf{x}$ can be written as $\mathbf{x}=uf(\mathbf{z})$, where $u$ is any word on any alphabet, $\mathbf{z}$ is a (necessarily binary) Sturmian sequence, and $f$ an aperiodic morphism from $\{0,1\}$ to any alphabet (see [28, 37, 72]). We state the following theorem.
Theorem 39.
Let $\mathbf{x}=yf(\mathbf{z})$ be a quasi-Sturmian sequence, where $y$ is any word, $\mathbf{z}$ is a Sturmian sequence, and $f$ is an aperiodic morphism from $\{0,1\}$ to any alphabet. Then
  • (a)
    either $f(\mathbf{z})$ is reversal-closed and $r_{\mathbf{x}}=\frac{n}{2}+O(1)$;
  • (b)
    or else $f(\mathbf{z})$ is not reversal-closed and $r_{\mathbf{x}}=n+O(1)$.
@@ 4.2 Quasi-Sturmian sequences
First we note that, clearly, $r_{\mathbf{x}}=r_{f(\mathbf{z})}(n)+O(1)$, hence it suffices to prove both statements for $r_{f(\mathbf{z})}$ instead of $r_{\mathbf{x}}$. Then we note that $\mathbf{z}$ and hence $f(\mathbf{z})$ are both uniformly recurrent. Thus we can apply Theorem 11 to $f(\mathbf{z})$. We now consider two cases.
  • (a)
    If $f(\mathbf{z})$ is reversal-closed, then, $r_{f(\mathbf{z})}=\frac{1}{2}(\rho_{f(\mathbf{z})}+\operatorname{Pal}_{f(\mathbf{z})})$. Since $f(\mathbf{z})$ is reversal-closed and quasi-Sturmian, we have from Theorem 17(b) that $\operatorname{Pal}_{f(\mathbf{z})}\leq 3$. Hence $r_{f(\mathbf{z})}(n)=\frac{1}{2}\rho_{f(\mathbf{z})}(n)+O(1)$ as desired.
  • (b)
    If $f(\mathbf{z})$ is not reversal-closed, then, $r_{f(\mathbf{z})}(n)=\rho_{f(\mathbf{z})}(n)$ for $n$ large enough. By assumption, we have $\rho_{f(\mathbf{z})}(n)=n+C$ for some constant $C$ and for $n$ large enough, so we have $r_{f(\mathbf{z})}(n)=n+O(1)$ as desired.
This ends the proof. ∎
Remark 40.
One can compare the first part of Theorem 39 above with Corollary 21.
4.3 Episturmian sequences
Among several generalizations of Sturmian sequences, episturmian sequences have in particular the property–sometimes even taken as part of their defini-tion–to be reversal-closed. Furthermore, their palindrome complexity has been studied. See the survey [50]; also see the survey [18]. We develop here for these sequences a theorem similar to Theorem 35 above.
@@ 4.3 Episturmian sequences
Let $A$ be a finite alphabet with cardinality $\ell$. A sequence $\mathbf{x}$ over $A$ is episturmian if it is reversal-closed and has at most one left special factor of each length. An episturmian sequence $\mathbf{x}$ is $\ell$-strict if it has exactly one left special factor of each length and for which every left special factor $u$ of $\mathbf{x}$ has $\ell$ distinct left extensions in $\mathbf{x}$.
@@
Let $\mathbf{x}$ be an $\ell$-strict episturmian sequence. Then, for all $n\geq 0$,
@@
Let $\mathbf{x}$ be an $\ell$-strict episturmian sequence. The case $n=0$ is true. Assume that $n\geq 1$. Then by [42, Theorem 7], we have $\rho_{\mathbf{x}}(n)=(\ell-1)n+1$. We also know that
@@
Using these together with Theorem 9(b), we deduce the desired result. ∎
@@
For the Tribonacci sequence $\mathbf{tr}$, which is the fixed point of the morphism $0\mapsto 01,1\mapsto 02,2\mapsto 0$, we have $r_{\mathbf{tr}}(n)=2\lfloor(n+1)/2\rfloor+1$ for all $n\geq 0$.
4.4 Billiard sequences on a hypercube
Since one interpretation of Sturmian sequences is the binary coding of irrational trajectories on a square billiard table, one can turn to irrational trajectories on a hypercube. The following result can be found in [16]: the first item is in [16, Corollary 1.6] and the second is the main theorem of that paper (which proves a conjecture due to Tamura; note the unexpected symmetry between $s$ and $n$, where $(s+1)$ is the dimension of the hypercube).
Theorem 44 ([16]).
Let $\mathbf{x}$ be an irrational billiard sequence on an $(s+1)$-dimensional hypercube.
  • (a)
    The sequence $\mathbf{x}$ is reversal-closed.
  • (b)
    The factor complexity of $\mathbf{x}$ is given by
    $\rho_{\mathbf{x}}(n)=\sum_{k=0}^{\min(s,n)}k!\binom{s}{k}\binom{n}{k}.$
Remark 45.
Note that, if $\mathbf{x}$ is an irrational billiard sequence on an $(s+1)$-dimensional hypercube, the previous result implies that $\rho_{\mathbf{x}}(n)=\Theta(n^{s})$. In particular for $s+1=2$, we obtain $\rho_{\mathbf{x}}(n)=n+1$ (which gives back Sturmian sequences), and for $s+1=3$, we obtain $\rho_{\mathbf{x}}(n)=n^{2}+n+1$, which had been conjectured by Rauzy and proved in [10].
Corollary 46.
Let $\mathbf{x}$ be an irrational billiard sequence on an hypercube of dimension $(s+1)$. Then its reflection complexity has the property that
$r_{\mathbf{x}}(n)\sim\frac{1}{2}\sum_{k=0}^{\min{(s,n)}}k!\binom{s}{k}\binom{n}{k}$
when $n$ tends to infinity. In particular, $r_{\mathbf{x}}(n)=\Theta(n^{s})$ when $n$ tends to infinity.
Proof.
Use Theorem 44 above with Corollary 21. ∎
4.5 Complementation-symmetric Rote sequences
So-called complementation-symmetric Rote sequences, which were defined and studied in [78], are related to Sturmian sequences as stated below in Theorem 48. In this section, after recalling their definition, we study their reflection complexity.
@@ 4.5 Complementation-symmetric Rote sequences
We consider the mapping $\Delta\colon\{0,1\}^{+}\to\{0,1\}^{*}$ defined as follows: $\Delta(a)=a$ for all $a\in\{0,1\}$ and for $n\geq 1$, $\Delta(v(0)v(1)\cdots v(n))$ $=$ $u(0)$ $u(1)$cdots $u(n-1)$ with $u(i)=(v(i+1)-v(i))\bmod{2}$ for all $i\in\{0,\ldots,n-1\}$. There is a natural extension of $\Delta$ to sequences: if $\mathbf{x}=(x(n))_{n\geq 0}$ is a binary sequence, then $\Delta(\mathbf{x})$ is the sequence whose $n$th letter is defined by $(x(n+1)-x(n))\bmod{2}$ for all $n\geq 0$. Observe that $\Delta(\mathbf{x})$ is the sequence of first differences of $\mathbf{x}$, taken modulo $2$.

### v4

@@ 3 General results
Let $\mathbf{x}$ be a primitive morphic sequence. We know that $\mathbf{x}$ is uniformly recurrent. Thus, from Theorem 11, either $\mathbf{x}$ is reversal-closed, or else it has no long palindromes. If $\mathbf{x}$ is reversal-closed, then by Corollary 22, we have $r_{\mathbf{x}}(n)\sim\frac{1}{2}\rho_{\mathbf{x}}(n)$. Otherwise, $\mathbf{x}$ has no long palindromes, then, still from Theorem 11, we have that $r_{\mathbf{x}}(n)=\rho_{\mathbf{x}}(n)$ for $n$ large enough. ∎
We conclude this section by leaving the following conjectures as open problems.
Conjecture 27.
Conjecture 28.
It can be shown that Conjecture 28 holds for the Thue–Morse, period-doubling, Golay–Shapiro, second-bit, paperfolding, Stewart choral, Baum-Sweet, Chacon, and Mephisto-Waltz sequences.
Conjecture 29.
4 Flye Sainte-Marie–Rauzy graphs for reflec-tion-equivalent classes
In this section, we prove the following inequality:
Theorem 30.
Let $\mathbf{x}$ be a sequence. Then $r_{\mathbf{x}}(n)\leq r_{\mathbf{x}}(n+1)+1$ for all $n\geq 0$.
Note that we can have equality in Theorem 30 for infinitely many values of $n$—see Example 16 above.
Initially, we discovered this growth property empirically. We prove it here by means of a graph construction related to the work of Flye Sainte-Marie. We remark that another recent application of this idea leads to a bound on the maximal number of powers in finite words [23, 64].
Flye Sainte-Marie graphs (also known as Rauzy graphs or De Bruijn graphs) for sequences $\mathbf{x}$ can be defined in the following way: for an integer $n\geq 0$, the graph $\Gamma_{\mathbf{x}}(n)$ is a directed graph whose vertex set is $\operatorname{Fac}_{\mathbf{x}}(n)$ and edge set is $\operatorname{Fac}_{\mathbf{x}}(n+1)$. An edge $e\in\operatorname{Fac}_{\mathbf{x}}(n+1)$ starts at the vertex $u$ and ends at the vertex $v$ if and only if $u$ is a prefix of $e$ and $v$ is a suffix. To simplify the notation, for words $w_{1},w_{2}$ of length $n$ and a word $w_{3}$ of length $n+1$, we write $w_{1}\xrightarrow{w_{3}}w_{2}$ if $w_{1}$ and $w_{2}$ are a prefix and a suffix of $w_{3}$, respectively.
Remark 31.
Flye Sainte-Marie first introduced the previous graph construction in [54] to represent the overlaps between all the binary words of the same length. De Bruijn rediscovered the same graph construction in [25] and acknowledged Flye Sainte-Marie’s priority of [54] later in [26]. Rauzy used the same construction in [81] to study the overlaps of factors in a word.
A directed graph is (weakly) connected if there is an undirected path between any pair of its vertices.
The following proposition is “folklore”. To the authors’ knowledge, the proof of this proposition does not seem to properly appear in the literature.
Proposition 32.
Let $\mathbf{x}$ be a sequence. For all $n\geq 0$, the graph $\Gamma_{\mathbf{x}}(n)$ is connected.
Proof.
Let $u$ and $v$ be two distinct length-$n$ factors of $\mathbf{x}$ and let us suppose that $u$ occurs before $v$ in $\mathbf{x}$. Then there are two integers $i<j$ such that $\mathbf{x}[i..i+n-1]=u$ and $\mathbf{x}[j..j+n-1]=v$. Then the path $\mathbf{x}[i..i+n],\mathbf{x}[i+1..i+n+1],\ldots,\mathbf{x}[j-1..j+n-1]$ connects $u$ and $v$. ∎
Remark 33.
Proposition 32 has already been (indirectly) mentioned in [85, p. 199, lines 4–6], which states that every graph $\Gamma_{\mathbf{x}}(n)$ necessarily has a vertex $o$ from which every other vertex can be reached by a direct path. Thus each pair of vertices $u$, $v$ can be connected by an undirected path passing through $o$.
Let $\mathbf{x}$ be a sequence. For all $n\geq 0$, let $\operatorname{ClassRef}_{\mathbf{x}}(n)$ be the set of all reflection-equivalent classes of length-$n$ factors of $\mathbf{x}$, i.e.,
$\operatorname{ClassRef}_{\mathbf{x}}(n)=\{\{u,u^{R}\}\,:\,u\in\operatorname{Fac}_{\mathbf{x}}(n)\},$
where $u$ may be equal to $u^{R}$. From the definition, the cardinality of the set $\operatorname{ClassRef}_{\mathbf{x}}(n)$ is $r_{\mathbf{x}}(n)$.
[…]
Remark 34.
From the definition, each element in $\operatorname{Fac}_{\mathbf{x}}(n+1)$ appears exactly once as an edge in $\Lambda_{\mathbf{x}}(n)$. This is easily seen by letting $e\in\operatorname{Fac}_{\mathbf{x}}(n+1)$ and by letting $u$ and $v$ be the length-$n$ prefix and suffix of $e$, respectively, and by using the property that $\{u,u^{R}\},\{v,v^{R}\}\in\operatorname{ClassRef}_{\mathbf{x}}(n)$. Thus, the number of vertices in $\Lambda_{\mathbf{x}}(n)$ is $r_{\mathbf{x}}(n)$ and the number of edges in $\Lambda_{\mathbf{x}}(n)$ is $\rho_{\mathbf{x}}(n+1)$.
Remark 35.
Letting $\mathbf{x}$ be a sequence, the graph $\Lambda_{\mathbf{x}}(n)$ is connected for arbitrary $n\geq 0$. This follows in a direct way from $\Gamma_{\mathbf{x}}(n)$ being connected from Proposition 32.
Lemma 36.
Let $\mathbf{x}$ be a sequence. Let $u,v$ be two distinct length-$(n+1)$ factors of $\mathbf{x}$ for some non-negative integer $n$. Then $u=v^{R}$ if and only if there are two distinct vertices $c_{1},c_{2}$ in $\Lambda_{\mathbf{x}}(n)$ such that $u$ is an edge from $c_{1}$ to $c_{2}$ and $v$ is an edge from $c_{2}$ to $c_{1}$.
Remark 37.
If there are two distinct elements in $c_{1}$ and/or $c_{2}$, there may be at most two distinct edges from $c_{1}$ to $c_{2}$ and vice-versa. So, to prove Lemma 36, we need to prove that for two distinct vertices, there is at most one edge from one to the other. Furthermore, if these two vertices are connected in both directions, then the two connecting edges in between are reflectively equivalent. Graphically, Lemma 36 gives us that $u=v^{R}$ and $u\neq v$ if and only if the graph $\Lambda_{\mathbf{x}}(n)$ has an occurrence of the pattern depicted in Figure 1.
$c_{1}$$c_{2}$$v$$u$
Figure 1: The consequence of Lemma 36 in the graph $\Lambda_{\mathbf{x}}(n)$, when $u=v^{R}$ but $u\not=v$.
Proof of Lemma 36.
If $u$ and $v=u^{R}$ are both edges of $\Lambda_{\text{{\bf x}}}(n)$, then we now show that $u=(c_{1},c_{2})$ and $v=(c_{2},c_{1})$ for some vertices $c_{1}$, $c_{2}$ of $\Lambda_{\text{{\bf x}}}(n)$ (see Fig. 1).
As indicated in Remark 37, we need to show that, in the case where $c_{1}\neq c_{2}$, there are at most two edges between two distinct vertices, and if there are exactly two edges, then one is the reverse of the other. In this direction, the given claim is clearly true for $n=0$ and $n=1$. For $n\geq 2$, we now show that if there exists $u$ such that $u\neq u^{R}$ and $u,u^{R}\in\operatorname{Fac}_{\mathbf{x}}(n+1)$, then there exist two distinct vertices satisfying the conditions in the statement. Let $u=au^{\prime}b$, where $a,b$ are letters. From the hypothesis that $u^{R}\in\operatorname{Fac}_{\mathbf{x}}(n+1)$, one has $au^{\prime},u^{\prime}b,u^{\prime R}a,bu^{\prime R}\in\operatorname{Fac}_{\mathbf{x}}(n)$. Thus, $c_{1}=\{au^{\prime},u^{\prime R}a\}$ and $c_{2}=\{u^{\prime}b,bu^{\prime R}\}$ are two vertices in the graph $\Lambda_{\mathbf{x}}(n)$. Moreover, since
$au^{\prime}\xrightarrow{au^{\prime}b}u^{\prime}b\text{ and }bu^{\prime R}\xrightarrow{bu^{\prime R}a}u^{\prime R}a,$
the word $au^{\prime}b$ is an edge from $c_{1}$ to $c_{2}$ and $bu^{\prime R}a$ is an edge from $c_{2}$ to $c_{1}$. Finally, to prove that $c_{1}\neq c_{2}$, we can use a proof by contradiction together with a case analysis by considering the $au^{\prime}=bu^{\prime R}$ case and the $au^{\prime}=u^{\prime}b$ case. A similar approach can be used to prove that $u=v^{R}$. We leave the details to the reader. ∎
Now we introduce the Flye Sainte-Marie graphs for reflection-equivalent classes. Let $\mathbf{x}$ be a sequence. For all $n\geq 0$, let $K_{\mathbf{x}}(n)$ be a directed graph such that its vertex set is $\operatorname{ClassRef}_{\mathbf{x}}(n)$ and its edge set is $\operatorname{ClassRef}_{\mathbf{x}}(n+1)$. Let $e\in\operatorname{ClassRef}_{\mathbf{x}}(n+1)$. We define the edges of $K_{\mathbf{x}}(n)$ as follows:
  • if $|e\cap\operatorname{Fac}_{\mathbf{x}}(n+1)|=1$, there exist a unique word $u\in e$ and two vertices $c_{1},c_{2}\in\operatorname{ClassRef}_{\mathbf{x}}(n)$ such that $u$ is from $c_{1}$ to $c_{2}$ in $\Lambda_{\mathbf{x}}(n)$. In the graph $K_{\mathbf{x}}(n)$, let $e$ be the edge from $c_{1}$ to $c_{2}$.
  • if $|e\cap\operatorname{Fac}_{\mathbf{x}}(n+1)|=2$, from Lemma 36 there exists a unique word $u\in e$ satisfying the following four conditions:
    • $u\neq u^{R}$;
    • $u$ is lexicographically smaller than $u^{R}$;
    • $u,u^{R}\in\operatorname{Fac}_{\mathbf{x}}(n+1)$;
    • there exist two vertices $c_{1},c_{2}\in\operatorname{ClassRef}_{\mathbf{x}}(n)$ such that $u$ is from $c_{1}$ to $c_{2}$ and $u^{R}$ is from $c_{2}$ to $c_{1}$ in $\Lambda_{\mathbf{x}}(n)$.
    In the graph $K_{\mathbf{x}}(n)$, let $e$ be the edge from $c_{1}$ to $c_{2}$.
Proposition 38.
Let $\mathbf{x}$ be a sequence. For all $n\geq 0$, the graph $K_{\mathbf{x}}(n)$ has exactly $r_{\mathbf{x}}(n)$ vertices and $r_{\mathbf{x}}(n+1)$ edges. Moreover, it is connected.
Proof.
From the construction, each element in $\operatorname{ClassRef}_{\mathbf{x}}(n+1)$ appears exactly once as an edge in $K_{\mathbf{x}}(n)$ and the vertex set of this graph is exactly $\operatorname{ClassRef}_{\mathbf{x}}(n)$. Thus, the number of vertices and edges is respectively $r_{\mathbf{x}}(n)$ and $r_{\mathbf{x}}(n+1)$. For the connectivity of the graph $K_{\mathbf{x}}(n)$, it is enough to show that every pair of vertices $c_{1},c_{2}$ connected by a single edge $u$ in $\Lambda_{\mathbf{x}}(n)$ is also connected by the single edge $\{u,u^{R}\}$ in $K_{\mathbf{x}}(n)$. Since the graph $\Lambda_{\mathbf{x}}(n)$ is connected, the graph $K_{\mathbf{x}}(n)$ is also connected. ∎
We can now complete the proof of Theorem 30:
Proof of Theorem 30.
This follows immediately from Proposition 38, since $K_{\mathbf{x}}(n)$ is a connected graph with $r_{\mathbf{x}}(n)$ vertices and $r_{\mathbf{x}}(n+1)$ edges. Thus $r_{\mathbf{x}}(n)\leq r_{\mathbf{x}}(n+1)+1$. ∎
@@ 5 Eventually periodic sequences
From Theorem 23 both sequences $(r_{\mathbf{x}}(2n))_{n\geq 0}$ and $(r_{\mathbf{x}}(2n+1))_{n\geq 0}$ are non-decreasing. Also, from the inequalities $\frac{1}{2}\rho_{\mathbf{x}}(n)\leq r_{\mathbf{x}}(n)\leq\rho_{\mathbf{x}}(n)$ in Theorem 9, and the fact that the sequence $(\rho_{\mathbf{x}}(n))_{n\geq 0}$ is non-decreasing, we have that either the three integer sequences $(r_{\mathbf{x}}(2n))_{n\geq 0}$, $(r_{\mathbf{x}}(2n+1))_{n\geq 0}$, and $(\rho_{\mathbf{x}}(n))_{n\geq 0}$ are all bounded, or else none of them is. Furthermore, we know that $(\rho_{\mathbf{x}}(n))_{n\geq 0}$ is bounded if and only if the sequence $\mathbf{x}$ is eventually periodic (Theorem 5 above). Hence, we have two cases depending on the periodicity of $\mathbf{x}$.
@@
    If $\mathbf{x}$ is not eventually periodic, its factor complexity is not bounded, thus both sequences $(r_{\mathbf{x}}(2n))_{n\geq 0}$ and $(r_{\mathbf{x}}(2n+1))_{n\geq 0}$ tend to infinity.
@@ 6 Sturmian sequences and generalizations
In this section, we study Sturmian sequences as well as some generalizations. First we state the following result, which characterizes Sturmian sequences in terms of their reflection complexity.
@@
Let $\mathbf{x}$ be a non-eventually periodic sequence over a finite alphabet.
@@
and $G_{n}$ be the subgroup of ${\mathcal{S}}_{n}$ generated by $\sigma_{n}$, i.e., the group $\{\sigma_{n},{\rm id}_{n}\}$. The number of distinct orbits of $\{1,2,\ldots,n\}$ under $G_{n}$ is equal to $n/2$ if $n$ is even, and to $(n+1)/2$ if $n$ is odd, which can be written $\lfloor(n+1)/2\rfloor$ in both cases. Thus, applying [36, Theorem 1] proves the first item of the theorem and the implication $\Longrightarrow$ of the second item.
To prove the last assertion, suppose that $\mathbf{x}$ is a Sturmian sequence. We know that every Sturmian sequence is reversal-closed (see [47, Theorem 4, p. 77], where reversals are called mirror images). Furthermore, it is proved in [47, Theorem 5, p. 77] that a sequence is Sturmian if and only if it has one palindrome of all even lengths and two palindromes of all odd lengths. Now, from Theorem 9(b) we have that
@@
This completes the proof. ∎
With regard to the above referenced work of Charlier et al. [36], also see the related and recent work by Luchinin and Puzynina [67].
@@
Recall that every quasi-Sturmian sequence $\mathbf{x}$ can be written as $\mathbf{x}=yf(\mathbf{z})$, where $y$ is a word over a finite alphabet, $\mathbf{z}$ is a (necessarily binary) Sturmian sequence, and $f$ an aperiodic morphism from $\{0,1\}$ to a finite alphabet, see [31, 40, 78]. (Recall that $f$ aperiodic means that $f(01)\neq f(10)$.) We state the following theorem.
Theorem 43.
Let $\mathbf{x}=yf(\mathbf{z})$ be a quasi-Sturmian sequence, where $y$ is a word, $\mathbf{z}$ is a Sturmian sequence, and $f$ is an aperiodic morphism from $\{0,1\}$ to a finite alphabet. Then
@@
    either $f(\mathbf{z})$ is reversal-closed and $r_{\mathbf{x}}(n)=\frac{n}{2}+O(1)$;
@@
    or else $f(\mathbf{z})$ is not reversal-closed and $r_{\mathbf{x}}(n)=n+O(1)$.
@@
Since $r_{\mathbf{x}}(n)=r_{f(\mathbf{z})}(n)+O(1)$, it suffices to prove both statements for $r_{f(\mathbf{z})}$ instead of $r_{\mathbf{x}}$. Since $\mathbf{z}$ and, hence, $f(\mathbf{z})$ are both uniformly recurrent, we can apply Theorem 11 to $f(\mathbf{z})$. The desired result then easily follows by considering the case whereby $f(\mathbf{z})$ is reversal-closed, so that $r_{f(\mathbf{z})}=\frac{1}{2}(\rho_{f(\mathbf{z})}+\operatorname{Pal}_{f(\mathbf{z})})$, and the case such that it is not reversal-closed, so that $r_{f(\mathbf{z})}(n)=\rho_{f(\mathbf{z})}(n)$. ∎
Among several generalizations of Sturmian sequences, episturmian sequences have in particular the property—sometimes even taken as part of their definition—to be reversal-closed. Furthermore, their palindrome complexity has been studied [18, 55]. Here we develop a theorem similar to Theorem 41 above for these sequences.
@@
Let $A$ be a finite alphabet with cardinality $\ell$. A sequence $\mathbf{x}$ over $A$ is episturmian if it is reversal-closed and has at most one left special factor of each length. An episturmian sequence $\mathbf{x}$ is $\ell$-strict if it has exactly one left special factor of each length and every left special factor $u$ of $\mathbf{x}$ has $\ell$ distinct left extensions in $\mathbf{x}$.
@@
Let $\mathbf{x}$ be an $\ell$-strict episturmian sequence. Then, for all $n\geq 0$, we have
@@
Let $\mathbf{x}$ be an $\ell$-strict episturmian sequence. The case $n=0$ is true. Assume that $n\geq 1$. Then by [46, Theorem 7], we have $\rho_{\mathbf{x}}(n)=(\ell-1)n+1$. We also know from [47] that
@@
Using these facts, together with Theorem 9(b), we deduce the desired result. ∎
@@
For the Tribonacci sequence $\mathbf{tr}$, which is the fixed point of the morphism $0\mapsto 01,1\mapsto 02,2\mapsto 0$, we have $r_{\mathbf{tr}}(n)=2\left\lfloor\frac{n+1}{2}\right\rfloor+1$ for all $n\geq 0$.
Since one interpretation of Sturmian sequences is the binary coding of irrational trajectories on a square billiard table, we turn our attention toward irrational trajectories on a hypercube. Using results due to Baryshnikov [16] together with Corollary 22, it follows in a direct way that
$r_{\mathbf{x}}(n)\sim\frac{1}{2}\sum_{k=0}^{\min{(s,n)}}k!\binom{s}{k}\binom{n}{k},$
when $n$ tends to infinity, for an irrational billiard sequence $\mathbf{x}$ on a hypercube of dimension $(s+1)$. In particular, we have that $r_{\mathbf{x}}(n)=\Theta(n^{s})$ when $n$ tends to infinity.
So-called complementation-symmetric Rote sequences, which were defined and studied by Rote in 1994 [85], are related to Sturmian sequences as stated below in Theorem 48.
@@
We consider the mapping $\Delta\colon\{0,1\}^{+}\to\{0,1\}^{*}$ defined as follows: $\Delta(a)=a$ for all $a\in\{0,1\}$ and
$\Delta(v(0)v(1)\cdots v(n))=u(0)u(1)\cdots u(n-1)$
for $n\geq 1$ with $u(i)=(v(i+1)-v(i))\bmod{2}$ for all $i\in\{0,\ldots,n-1\}$. There is a natural extension of $\Delta$ to sequences: if $\mathbf{x}=(x(n))_{n\geq 0}$ is a binary sequence, then $\Delta(\mathbf{x})$ is the sequence whose $n$th letter is defined by
$(x(n+1)-x(n))\bmod{2}$
for all $n\geq 0$. Observe that $\Delta(\mathbf{x})$ is the sequence of first differences of $\mathbf{x}$, taken modulo $2$.